OpenStudy (joshawad):

help i forgot how to do algebraic fractions! i knoiw its kind of stupid but: 1/2B+1/3=2/9

OpenStudy (anonymous):

Alright, do you know how to isolate the variable?

OpenStudy (joshawad):

no i forgot...

OpenStudy (joshawad):

typing a book are you? lol

OpenStudy (anonymous):

Okay. So, you see where it says 1/2 B? 1/2 is the same as 1 divided by 2, so we divide 1 by 2, getting 0.5. Then, the B is attached to the 0.5, and we need to make the B not attached to anything. So, because it is 0.5 TIMES B, we need to DIVIDE each side of the equals sign by 0.5. Well, when you are isolating a variable, you divide one side of the equals sign by 0.5 and instead of dividing 0.5 by 0.5, they just cancel each other out. From that, you get B + 1/3 = 0.44. Then, we need to subtract 1/3 from each side of the equals sign, and we get B = 0.111. What are your possible answers, because there is a way to change this into a fraction if need be. Sorry for taking so long- I had to do it on pencil and paper first. :)

OpenStudy (texaschic101):

If you want to get rid of the fractions, multiply the entire equation by the LCD, which is 18. This will leave you with just whole numbers

OpenStudy (joshawad):

ok thanks can i was gonna make a new thread but how about .03Y+1=4.3

OpenStudy (joshawad):

how u get 18

OpenStudy (anonymous):

The LCD is the Lowest Common Denominator. The denominator is the bottom, like the 3 in 1/3 or the 6 in 4/6. Take is away, @texaschic101

OpenStudy (anonymous):

*It

OpenStudy (texaschic101):

sorry about that...I lost connection

OpenStudy (texaschic101):

1/2b + 1/3 = 2/9 (multiply by the LCD which is 18) 9b + 6 = 4 -- subtract 6 from both sides 9b = 4 - 6 9b = 2 b = 2/9 .03y + 1 = 4.3 .03y = 4.3 - 1 .03y = 3.3 .y = 3.3/.03 y = 110

OpenStudy (texaschic101):

oops...made a mistake on the first one..... 9b = 4 - 6 9b = -2 b = -2/9

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (joshawad):

yes how did you get rid of the denominator of 18 my mom grades my work and shes confused as well

OpenStudy (texaschic101):

18(1/2b) + 18(1/3) = 18(2/9) 18/2b + 18/3 = 36/9 9b + 6 = 4 9b = 4 - 6 9b = -2 b = -2/9

OpenStudy (joshawad):

you just mutiplied everything by 18?

OpenStudy (texaschic101):

if you have a bunch of fractions and you do not want to deal with the fractions, multiply the entire equation by the LCD and it will get rid of the fractions and turn them into whole numbers

OpenStudy (joshawad):

Ok thanks!

OpenStudy (texaschic101):

no problem :)